# The Essential Role of Causality in Foundation World Models for Embodied AI

### v1

@@ Abstract
Recent advances in foundation models, especially in large multi-modal models and conversational agents, have ignited interest in the potential of generally capable embodied agents. Such agents would require the ability to perform new tasks in many different real-world environments. However, current foundation models fail to accurately model physical interactions with the real world thus not sufficient for Embodied AI. The study of causality lends itself to the construction of veridical world models, which are crucial for accurately predicting the outcomes of possible interactions. This paper focuses on the prospects of building foundation world models for the upcoming generation of embodied agents and presents a novel viewpoint on the significance of causality within these. We posit that integrating causal considerations is vital to facilitate meaningful physical interactions with the world. Finally, we demystify misconceptions about causality in this context and present our outlook for future research.
@@ 1 Introduction
The ability to perform physically meaningful interactions in a real-world environment is essential to maximize the potential of technology in everyday life. Entities capable of conducting interactions of this kind are in our work referred to as embodied agents. These agents can vary in form, ranging from robots to humans augmented by AI in mixed reality environments, or in virtual reality where interactions bear physical importance.
Foundation models have recently accounted for the most significant advancement in generalization across a range of tasks with multi-modal input, compared to canonical deep learning models from the past decade. In particular, a single model learns to address a wide variety of tasks in a zero-shot manner. In robotics, we anticipate that foundation models for embodied agents will make the introduction of new robots into our life effortless: each new robot can immediately solve a given task using the same model without the need for a specialized model. In addition, robots can quickly adapt to a new environment by interacting with it and use the new experiences to improve without lengthy retraining or complex programming.
@@
Causality at its core aims to understand the consequences of actions, allowing for interaction planning. Philosophy and cognitive sciences sees understanding causal concepts as both the basics and sometimes the ultimate goal for humans to learn how to interact with the world (Gibson, 1978; Gopnik et al., 2007; Adams & Aizawa, 2021) and key in children’s development (Piaget, 1965; Gibson, 1988). This aligns with the concept of affordances in Embodied AI research (Kjellström et al., 2011; Koppula et al., 2013; Koppula & Saxena, 2013; Ardón et al., 2020; Ahn et al., 2022), which pertains to the types of action that can be applied and the resulting consequences. Importantly, even with the help of available real or simulated environments, the experimentation might still be too coarse-grained to deal with spurious relationships (Herd & Miles, 2019; Lavin et al., 2021). Thus, the world model for an agent, including humans, should be causal-aware (Anonymous, 2024b) and able to efficiently incorporate data from demonstration and interaction (i.e., the property referred as veridicality in this paper). As such, causality offers tools and insights that hold the key pieces to building Foundation Veridical World Models (FVWMs) (Section 2.2) that will power future embodied agents.
In this position paper, we argue that future progress in Embodied AI relies on advances in foundation models, given the readiness of various technological considerations (Section 2). For such research, causality is pivotal in developing the FVWM, which requires us to go beyond canonical causal research paradigms (Section 3). Thus, we foresee research opportunities (Section 4) advancing foundation world models in Embodied AI from a causal perspective. We extend our discussion with considerations in deploying Embodied AI (Section 5) in the end.
@@ 2 Embodied AI and Foundation Veridical World Models
Figure 1: The foundation world model is designed to comprehend associations, counterfactuals, and interactions within the world, aiding any embodied AI agent in executing tasks across diverse environments and platforms.
For many years, research has pursued the creation of real Embodied AI. Yet, the tangible success of these technologies is predominantly seen in scenarios where environments are tailored to suit the robot for a particular task, such as assembling machines in production lines. The inability to adapt to diverse environments and multiple tasks limits their real-world applicability and the range of problems it can solve. It highlights a significant gap in fulfilling the promises of embodied AI to be a multi-task solver in general environments. We believe that recent technological advancements (Section 2.1), including hardware, foundation models, and data, have brought general Embodied AI within reach, where the remaining gap can be overcome by veridical world models (Section 2.2).
@@ 2.1 The readiness of Embodied AI research
The pace of innovation in hardware and systems, ranging from real-world robots to virtual and mixed reality devices, is advancing at an unprecedented rate, though there is still a significant journey ahead before achieving maturity for general tasks. The costs of both humanoid (e.g. Optimus, Phoenix (Sanctuary, 2024), Digit (Agility, 2024)) and quadrupedal robots (e.g. Spot (BostonDynamics, 2024), Go2 (Unitree, 2024) ) have been rapidly decreasing over the last decade, with a huge potential market size (GoldmanSachs, 2023; Tilley, 2017). For virtual mixed reality experiences, Oculus from Meta has increased in popularity and Apple’s Vision Pro is taking a commercial leap offering new interactive experiences, creating opportunities for virtual agents. In addition, new types of sensors for robots have been invented recently. Electronic skin can capture the sensation of touch (Sanderson, 2021) to enable dexterous robots (Lambeta et al., 2020; Sun et al., 2022), while perovskite retinomorphic sensors (Trujillo Herrera & Labram, 2020) mimic the human retina and react to changes in illumination rather than constant signals. These advancements offer different ways to sense the world. This enables the consideration of more challenging manipulations, such as those involving deformable objects, rather than rigid ones (Yin et al., 2021).
Meanwhile, the evolution of AI computing power has accelerated at an unprecedented rate. This surge in computing capabilities is primarily driven by innovations in hardware, such as more powerful GPUs and specialized AI processors (Dally et al., 2021; Saravanan & Kouzani, 2023), providing the computing basis for foundation models. These developments have enabled AI systems to process and analyze vast amounts of data more quickly and more effectively than ever before, thereby expanding the horizons of potential across various domains.
@@
The paradigm shift driven by foundation models in both language and vision has markedly enhanced the AI’s ability to understand and perceive the environment (Bubeck et al., 2023; Achiam et al., 2023; Yuan et al., 2021; Liu et al., 2023; Chen et al., 2023). These systems now demonstrate an increasing capability to generalize on high-level planning and decision-making tasks, often in a zero-shot manner. This has led to a shift in the research paradigm from specialized models for a single target task and dataset to foundation models generalizing across tasks and multiple datasets. The increased generalization capabilities hold the key to build foundation veridical world models required for embodied AI, opening the doors to applications that were once confined to the realms of imagination.
@@
Adding to the timeliness of this venture is the increasing amount of of data in the field of embodied AI, which is crucial for effective training of foundation world models. Robotics simulation environments and demonstration datasets have become more diverse, encompassing different types of tasks and scenarios, such as OpenX embodiment (Open X-Embodiment Collaboration et al., 2023), Isaac gym (Makoviychuk et al., 2021; Ma et al., 2023), CloudGripper (Zahid & Pokorny, 2023), Habitat 3.0 (Puig et al., 2023) and Causal World simulations (Ahmed et al., 2020; Bauer et al., 2022), greatly increasing the amount of available data. In addition, the maturity of physical robots and sensors also facilitates the generation of diverse multi-modal data. These real-world datasets represent different aspects of the environment. Moreover, sources like YouTube provide a wealth of information in the form of videos or general human behaviors.
@@ 2.2 Towards Foundation Veridical World Models
Figure 2: A veridical world model can predict the consequences of sequence of actions and corresponding counterfactual correctly with confidence, ignoring irrelevant information like color of the cup. For example, the model should correctly predict that the cup will break if ‘open drawer’ action is not taken in the sequence. Images taken from AI2THOR (Kolve et al., 2017) simulator.
We believe that world models hold great promise for the future of embodied AI. One ideal world model is to facilitate the entire pipeline, spanning various tasks and platforms as shown in Figure 1. To achieve that, a desirable property is the ability to consume multi-modal inputs, generalize across environments, domains and tasks (i.e., being foundational). This requires the model to faithfully understand and model the world dynamics, i.e., being veridical. It needs to go beyond merely summarizing the content given what is observed, but also truly comprehend the world in a way that enables actionable interventions.
Combining the above two key aspects leads to the idea of FVWMs, or simply Foundation World Models (FWMs): a single (or set of) veridical, multi-modal models that can:
@@
    (Foundational) Being able to generalize (i) and (ii) across diverse systems or domains encountered in the world.
When only (i) and (ii) are satisfied, FVWMs reduces to world models, originally intended to simulate possible futures of a single system (Ha & Schmidhuber, 2018; Li et al., 2020). By definition, FVWMs require a flexible formulation that can, using observational and experimental data from the world, reveal the underlying principles. This naturally falls into the realm of causality. However, incorporating such reasoning abilities while being able to process and generalize across multi-modal datasets and tasks is challenging, indicating substantial research opportunities for significant real-world impact.
@@ 3.1 Canonical approaches to causal modelling
Given the difficulty to access the true system, theory-oriented research prevails in causality. For example, identifiability theory concerns the theoretical feasibility of deducing causal relations, interventional or counterfactual effects from observational data, assuming a certain set of models. Once identifiability is determined, the next steps involve developing methods for identifying (Pearl, 2009a; Angrist et al., 1996) and accurately estimating these causal quantities (Rosenbaum & Rubin, 1983; Stuart, 2010; Li et al., 2018; Chernozhukov et al., 2018; Wager & Athey, 2018). This theory-driven approach is one way to ensure the precision and trustworthiness of the model, as it contributes significantly to our comprehension of underlying causal mechanisms and guides decision making in relevant domains. However, this approach may also limit the method development considering real-world impact (Imbens, 2020).
@@
The canonical theory-oriented research has facilitated significant progress. Meanwhile, we are at a historical juncture to rethink this strategy for FVWMs and embodied AI due to the following limitations. Firstly, real-world sensory inputs are usually not structured into well-defined variables that admit a causal explanation (in terms of SEM/PO) and meaningful interventions (Schölkopf et al., 2021). Secondly, in many real-world scenarios, certain key theoretical properties, such as identifiability, may not hold, and their associated assumptions are often not verifiable (Pearl, 2009b; Robins, 2000). Thirdly, in the context of FVWMs, the case-specific nature of causal identification complicates the integration of causal reasoning (Bareinboim et al., 2022). Foundation models aim to provide a generalizable understanding of the world, but the bespoke nature of causal identification, which often relies on domain-specific assumptions, poses a significant barrier for the general environment. This discordance reveals a gap between the theoretical guarantee of causal reasoning and the practical needs of causal principles in foundation models, necessitating a re-evaluation of the approaches used to embed causality for FVWMs and embodied AI.
@@ 3.2 Demystifying causal modeling for FVWMs
The canonical causal research paradigm has inadvertently contributed to certain misconceptions regarding the field of causal machine learning, thus limiting its broader impact, especially under FVWMs and embodied AI. In the following, we aim to demystify three main misconceptions: (1) casual research is a separate field from Embodied AI and FVWMs; (2) causal research must be theory-oriented; and (3) predictive and causal models are different.
@@
Despite causality being pivotal for the generalization of world models (Anonymous, 2024b), Causal ML is often overlooked in discussions, compared to RL and foundation models. This oversight partly stems from distinct tools, such as SEMs and PO, and application domains traditionally associated with canonical causal research. However, it is crucial to recognize that Causal ML is intricately connected to both reinforcement learning (bandit and imitation learning as a special case) and foundation models with shared goals.
@@
Bandits and RL share the same goal of learning the best action policy from observational data. Offline RL fundamentally shares the same goal as causal modeling, while online RL focuses more on learning from interactions (Kaddour et al., 2022). While RL does not always explicitly model environmental dynamics – as seen in model-free RL – it inherently captures these dynamics within the learned policy or value function through interactions with environment. Thus, RL is intrinsically causal in a board sense considering that it need to inherently understand the consequence of actions for policy learning. A more explicit connection can be seen in model-based RL, where temporal causal models can be used to formulate the transition dynamics. Such temporal causal models induce the transition probabilities between states across time. Actions by agents are then viewed as interventions to this model. Leveraging ideas from causality would allow the learned dynamics to only reason over causally relevant attributes such as force and location to predict the next state (e.g. location of the cup in Figure 2). Many recent works have begun to incorporate causal concepts into the state transition dynamics (see Section 6). This potentially enhances sample efficiency and generalizability beyond training contexts.
@@
Large multi-modal foundation models aim to facilitate reasoning and planning (Sun et al., 2023; Huang & Chang, 2022; Webb et al., 2023; Li et al., 2023), which inherently requires understanding the physical nature of environments. Although much progress has been made (Wei et al., 2022; Zhang et al., 2022; Creswell et al., 2022; Wang et al., 2023a; Green et al., 2012), they are trained to gather information only at the observational level, which may limit their understanding of the physical world (lack of interventional and counterfactual information). While these models, including CLIP (Radford et al., 2021) and GPT-4 (Achiam et al., 2023), excel in various tasks like vision-language understanding and image recognition with multi-modal input, they are limited by issues like hallucination (Huang et al., 2023a; Gunjal et al., 2023) and reduced multi-step reasoning abilities (Jin et al., 2024; Sun et al., 2023), indicating the lack of causal understanding. On the other hand, the canonical causality research fails to incorporate multi-modal dataset and is typically tied with predefined variables for SEM. Recent work (discussed in Section 6) reveals the synergy between Causal ML and foundation models for mutual enhancement, and paves the way for FVWMs.
@@
The traditional emphasis of theory-oriented approach in causal machine learning research often overshadows the empirical-driven approach, that prioritizes the causal experimentation, verification, and hypothesis selection (Dorie et al., 2019; Jensen, 2019; Gong et al., 2022; Nie & Wager, 2021; Saito & Yasui, 2020; Vansteelandt et al., 2012; Schuler et al., 2018; Rolling & Yang, 2014; Matthieu & Gaël, 2023; Cook et al., 2002; Hoover, 2001; Bareinboim & Pearl, 2016; Bareinboim, 2014). Consequently, over-reliance on the pure theoretical approach can lead to models that fail to perform effectively in practical, complex scenarios. We would like to emphasize that one of the end goals of causality research is causal-aware reasoning and to predict the consequences accurately given any actions. Canonical formulations with theoretical guarantee is one way for causal research, while focusing on evaluating the causal property and empirical results is an alternative when it comes to FVWMs. Such an empirical-driven shift allows models to assimilate causal knowledge from a wide array of scenarios, akin to the way large language models are evaluated. Moreover, the unique problem setting of embodied AI opens up a realm where empirical evaluation can become the cornerstone of causal model assessment. Through interaction with both real and simulated environments, embodied AI agents can generate and leverage interventional or counterfactual data, thereby permitting a more pragmatic approach to causal understanding. Thus, the empirical-driven causal research under FVWMs and embodied AI is equally or more important than the theory-oriented paradigm.
@@
The commonly understood principle that correlation does not imply causation is particularly relevant in the field of predictive modeling, and it is often misunderstood as predictive models cannot be causal. This interpretation holds some truth when models are trained solely on observational data, which are prone to confounding effects that can bias predictions about the outcomes of specific actions. However, this changes with access to active intervention data, allowing the predictive model to demonstrate causal properties. This aligns with the discussion in misconception I : RL models are inherently causal and can learn to take optimal actions for a variety of tasks (Kaddour et al., 2022; Anonymous, 2024b). Thus, causal models are inherently high-accuracy prediction models of consequences about all possible actions. Predictive models can also be considered causal models when trained with appropriate algorithms and datasets, even without explicit causal theory considerations. In the realm of foundation models for embodied AI, greater opportunities arise from the potential interaction with the physical world. This allows the integration of both causal considerations and large-scale prediction models (Ke et al., 2019; Scherrer et al., 2021; Toth et al., 2022).

### v2

@@ Abstract
Recent advances in foundation models, especially in large multi-modal models and conversational agents, have ignited interest in the potential of generally capable embodied agents. Such agents will require the ability to perform new tasks in many different real-world environments. However, current foundation models fail to accurately model physical interactions and are therefore insufficient for Embodied AI. The study of causality lends itself to the construction of veridical world models, which are crucial for accurately predicting the outcomes of possible interactions. This paper focuses on the prospects of building foundation world models for the upcoming generation of embodied agents and presents a novel viewpoint on the significance of causality within these. We posit that integrating causal considerations is vital to facilitating meaningful physical interactions with the world. Finally, we demystify misconceptions about causality in this context and present our outlook for future research.
@@ 1 Introduction
The ability to perform physically meaningful interactions in real-world environments is essential to maximizing the potential of technology in everyday life. Entities capable of conducting interactions of this kind are in our work referred to as embodied agents. These agents can vary in form, ranging from mechanical robots to humans augmented by AI in mixed or virtual reality environments where interactions bear physical importance.
Foundation models account for a significant advancement in generalization across a range of tasks with multi-modal input, compared to canonical deep learning models from the past decade. A single foundation model can address a wide variety of tasks in a zero-shot manner. For example, in robotics, we anticipate that foundation models for embodied agents will make the introduction of new robots into our life effortless: each new robot can immediately solve a given task using the same model without the need for a specialized model. In addition, robots can quickly adapt to a new environment by interacting with it and using the new experiences to improve without lengthy retraining or complex programming.
@@
World models (Ha & Schmidhuber, 2018) make action-conditioned predictions about the future state of the environment by training on action-labelled interaction data. World models therefore capture the consequences of actions and are suitable for optimising decision-making (Sutton, 1991; Janner et al., 2019). To train general world models it will be necessary to leverage large readily-available observational datasets, such as internet videos, in addition to action-labelled data which is much more scarce and expensive to generate. The best way to utilise observational data for world modeling remains an open question (Seo et al., 2022; Bruce et al., 2024). In this paper we argue that causal considerations are necessary to unlock the potential for training world models from internet-scale data.
Causality at its core aims to understand the consequences of actions, allowing for interaction planning. Philosophy and cognitive sciences sees understanding causal concepts as both the basics and sometimes the ultimate goal for humans to learn how to interact with the world (Gibson, 1978; Gopnik et al., 2007; Adams & Aizawa, 2021) and key in children’s development (Piaget, 1965; Gibson, 1988). This aligns with the concept of affordances in Embodied AI research (Kjellström et al., 2011; Koppula et al., 2013; Koppula & Saxena, 2013; Ardón et al., 2020; Ahn et al., 2022), which pertains to the types of action that can be applied and the resulting consequences. Importantly, even with the help of available real or simulated environments, the experimentation might still be too coarse-grained to deal with spurious relationships (Herd & Miles, 2019; Lavin et al., 2021). Thus, the world model for an agent, including humans, should be causal-aware (Richens & Everitt, 2024) and able to efficiently incorporate data from observations, demonstrations, and interaction (i.e., the property referred as veridicality in this paper). As such, causality offers tools and insights that hold the key pieces to building Foundation Veridical World Models (FVWMs) (Section 2.2) that will power future embodied agents.
In this position paper, we argue that future progress in Embodied AI relies on further advances in foundation models (Section 2). For such research, causality is pivotal in developing the FVWM, which requires us to go beyond canonical causal research paradigms (Section 3). Thus, we foresee research opportunities (Section 4) advancing foundation world models in Embodied AI from a causal perspective. We extend our discussion with considerations for the eventual deployment of Embodied AI (Section 5).
@@ 2 Embodied AI and Foundation Veridical World Models
Figure 1: The Foundation Veridical World Model (FVWM) is designed to comprehend associations, counterfactuals, and interactions within the world, aiding any embodied AI agent in executing tasks across diverse environments and platforms.
For many years, research has pursued the creation of Embodied AI. However, existing approaches are limited to performing narrow tasks in highly constrained environments, such as assembling machines in production lines. This inhibits the applicability of existing Embodied AI systems to general real-world settings, where environments are often dynamic and unstructured, and the set of tasks to be performed is open-ended. This highlights a significant gap between current technologies and fulfilling the promises of general Embodied AI. We believe that recent technological advancements (Section 2.1), including hardware, foundation models, and data, have brought general Embodied AI within reach, where the remaining gap can be overcome by veridical world models (Section 2.2).
@@ 2.1 The readiness of Embodied AI research
The pace of innovation in hardware and systems, ranging from real-world robots to virtual and mixed reality devices, is advancing at an unprecedented rate, though there is still a significant journey ahead before achieving maturity for general tasks. The costs of both humanoid (e.g. Optimus, Phoenix (Sanctuary, 2024), Digit (Agility, 2024)) and quadrupedal robots (e.g. Spot (BostonDynamics, 2024), Go2 (Unitree, 2024) ) have been rapidly decreasing over the last decade, with a huge potential market size (GoldmanSachs, 2023; Tilley, 2017). For virtual mixed reality experiences, Oculus from Meta has increased in popularity and Apple’s Vision Pro is taking a commercial leap into offering new interactive experiences, creating opportunities for virtual agents. In addition, new types of sensors for robots have been invented recently. Electronic skin can capture the sensation of touch (Sanderson, 2021) to enable dexterous robots (Lambeta et al., 2020; Sun et al., 2022), while perovskite retinomorphic sensors (Trujillo Herrera & Labram, 2020) mimic the human retina and react to changes in illumination rather than constant signals. These advancements offer different ways to sense the world. This enables the consideration of more dextrous manipulation tasks, such as those involving deformable objects, rather than rigid ones (Yin et al., 2021).
Meanwhile, the evolution of AI computing power has accelerated at an unprecedented rate. This surge in computing capabilities is primarily driven by innovations in hardware, such as more powerful GPUs and specialized AI processors (Dally et al., 2021; Saravanan & Kouzani, 2023), providing the computing basis for foundation models. These developments have enabled AI systems to process vast amounts of data more quickly and more effectively than ever before, expanding the horizons of potential across various domains.
@@
The paradigm shift driven by foundation models in both language and vision has markedly enhanced the ability of AI models to understand and perceive the environment (Bubeck et al., 2023; Achiam et al., 2023; Yuan et al., 2021; Liu et al., 2023; Chen et al., 2023). These systems now demonstrate an increasing capability to generalize on high-level planning and decision-making tasks, often in a zero-shot manner (Ahn et al., 2022). This has led to a shift in the research paradigm from specialized models for a single target task and dataset to foundation models generalizing across many tasks and datasets. These increased generalization capabilities hold the key to building world models suitable for general Embodied AI systems.
@@
Adding to the timeliness of this venture is the increasing amount of of data in the field of Embodied AI, which is crucial for effective training of foundation world models. Robotics simulation environments and demonstration datasets have become more diverse, encompassing different types of tasks and scenarios, such as OpenX embodiment (Open X-Embodiment Collaboration et al., 2023), Isaac gym (Makoviychuk et al., 2021; Ma et al., 2023), CloudGripper (Zahid & Pokorny, 2023), Habitat 3.0 (Puig et al., 2023) and Causal World simulations (Ahmed et al., 2020; Bauer et al., 2022), greatly increasing the amount of available data. In addition, the maturity of physical robots and sensors also facilitates the generation of diverse multi-modal data. These real-world datasets represent different aspects of the environment. Moreover, sources like YouTube provide a wealth of information in the form of videos or general human behaviors (Seo et al., 2022).
@@ 2.2 Towards Foundation Veridical World Models
Figure 2: A veridical world model can predict the consequences of a sequence of actions and corresponding counterfactuals correctly with confidence, ignoring irrelevant information like color of the cup. For example, the model should correctly predict that the cup will break if ‘open drawer’ action is not taken in the sequence. Images taken from AI2THOR (Kolve et al., 2017) simulator.
We believe that world models hold great promise for the future of Embodied AI. Ideally, a general world model will facilitate the entire Embodied AI pipeline, spanning various tasks and platforms as shown in Figure 1. To achieve that, a required property is the ability to consume multi-modal inputs, and generalize across environments, domains, and tasks (i.e., being foundational). This requires the model to faithfully understand and model the world dynamics, i.e., being veridical. It needs to go beyond merely summarizing the content given what is observed, but also truly comprehend the world in a way that enables actionable interventions.
Combining the above two key aspects leads to the idea of FVWMs, or simply Foundation World Models (FWMs), a single (or set of) veridical, multi-modal models that can:
@@
    (Foundational) Generalize (i) and (ii) across diverse systems or domains encountered in the world.
When only (i) and (ii) are satisfied, FVWMs reduces to world models, originally intended to simulate possible futures of a single system (Ha & Schmidhuber, 2018; Li et al., 2020). To accomplish (iii), FVWMs require a flexible formulation that can, using observational and interaction data from diverse real-world datasets, reveal the underlying principles that generalize across environments. This naturally falls into the realm of causality. However, incorporating such causal reasoning abilities while being able to process and generalize across multi-modal datasets and tasks is challenging, indicating substantial research opportunities for significant real-world impact.
@@ 3.1 Canonical approaches to causal modelling
Given the difficulty of accessing the true system, theory-oriented research prevails in causality. For example, identifiability theory concerns the theoretical feasibility of deducing causal relations, interventional or counterfactual effects from observational data, assuming a certain set of models. Once identifiability is determined, the next steps involve developing methods for identifying (Pearl, 2009a; Angrist et al., 1996) and accurately estimating these causal quantities (Rosenbaum & Rubin, 1983; Stuart, 2010; Li et al., 2018; Chernozhukov et al., 2018; Wager & Athey, 2018). This theory-driven approach is one way to ensure the precision and trustworthiness of the model, as it contributes significantly to our comprehension of underlying causal mechanisms and guides decision making in relevant domains. However, as we shall discuss, this approach may limit the real-world impact of methods developed (Imbens, 2020).
@@
The canonical theory-oriented research has facilitated significant progress. Meanwhile, we are at a historical juncture to rethink this strategy for FVWMs and embodied AI due to the following limitations. First, real-world sensory inputs are usually not structured into well-defined variables that admit a causal explanation (in terms of SEM/PO) and meaningful interventions (Schölkopf et al., 2021). Second, in many real-world scenarios, certain key theoretical properties, such as identifiability, may not hold, and their associated assumptions are often not verifiable (Pearl, 2009b; Robins, 2000). Third, in the context of FVWMs, the case-specific nature of causal identification complicates the integration of causal reasoning (Bareinboim et al., 2022). Foundation models aim to provide a generalizable understanding of the world, but the bespoke nature of causal identification, which often relies on domain-specific assumptions, poses a significant barrier for general environments. This discordance reveals a gap between the theoretical guarantee of causal reasoning and the practical needs of causal principles in foundation models, necessitating a re-evaluation of the approaches used to embed causality for FVWMs and embodied AI.
@@ 3.2 Demystifying causal modeling for FVWMs
The canonical causal research paradigm has inadvertently contributed to certain misconceptions regarding the field of causal machine learning, thus limiting its broader impact, especially in the context of FVWMs and embodied AI. In the following, we aim to demystify three main misconceptions: (1) casual research is a separate field from Embodied AI and FVWMs; (2) causal research must be theory-oriented; and (3) predictive and causal models are different.
@@
Despite causality being pivotal for the generalization of world models (Richens & Everitt, 2024), Causal ML is often overlooked in Embodied AI discussions, compared to RL and foundation models. This oversight partly stems from distinct tools, such as SEMs and PO, and the application domains traditionally associated with canonical causal research. However, it is crucial to recognize that Causal ML is intricately connected to both reinforcement learning (bandits and imitation learning as a special case) and foundation models, with shared goals.
@@
Bandits and online RL share the aim of learning the best action policy from online interactions. Offline RL assumes access to only fixed dataset from the system of interest, and therefore considers the same problem setting as causal modeling (Kaddour et al., 2022). While RL does not always explicitly model environmental dynamics – as seen in model-free RL – it implicitly captures these dynamics within the learned policy or value function. Thus, RL is intrinsically causal in the sense that it needs to inherently understand the consequences of actions for policy learning. A more explicit connection can be seen in model-based RL, where temporal causal models can be used to formulate the transition dynamics. Such temporal causal models induce the transition probabilities between states across time. Actions by agents are then viewed as interventions to this model. Leveraging ideas from causality would allow the learned dynamics to only reason over causally relevant attributes such as force and location to predict the next state (e.g. location of the cup in Figure 2). Many recent works have begun to incorporate causal concepts into the state transition dynamics (see Section 6). This potentially enhances sample efficiency and generalizability beyond training contexts.
@@
Large multi-modal foundation models aim to facilitate reasoning and planning (Sun et al., 2023; Huang & Chang, 2022; Webb et al., 2023; Li et al., 2023), which inherently requires understanding the physical nature of environments. Although much progress has been made (Wei et al., 2022; Zhang et al., 2022; Creswell et al., 2022; Wang et al., 2023a; Green et al., 2012), they are trained to gather information only at the observational level, which may limit their understanding of the physical world (lack of interventional and counterfactual information). While these models, including CLIP (Radford et al., 2021) and GPT-4 (Achiam et al., 2023), excel in various tasks like vision-language understanding and image recognition with multi-modal input, they are limited by issues like hallucination (Huang et al., 2023a; Gunjal et al., 2023) and reduced multi-step reasoning abilities (Jin et al., 2024; Sun et al., 2023), indicating the lack of causal understanding. On the other hand, canonical causality research fails to incorporate multi-modal datasets and is typically tied with predefined variables for SEM. Recent work (discussed in Section 6) reveals the synergy between Causal ML and foundation models for mutual enhancement, and paves the way for FVWMs.
@@
The traditional emphasis on theory-oriented approaches in causal machine learning research often overshadows empirically-driven approaches, which prioritize causal experimentation, verification, and hypothesis selection (Dorie et al., 2019; Jensen, 2019; Gong et al., 2022; Nie & Wager, 2021; Saito & Yasui, 2020; Vansteelandt et al., 2012; Schuler et al., 2018; Rolling & Yang, 2014; Matthieu & Gaël, 2023; Cook et al., 2002; Hoover, 2001; Bareinboim & Pearl, 2016; Bareinboim, 2014). Consequently, over-reliance on purely theoretical approaches can lead to models that fail to perform effectively in practical, complex scenarios. We would like to emphasize that one of the end goals of causality research is causal-aware reasoning and the ability to predict the consequences of any action accurately. Canonical formulations with theoretical guarantees are one approach to causal research, while focusing on empirical evaluations of causal properties is an alternative when it comes to FVWMs. A shift towards empirically-driven research allows models to assimilate causal knowledge from a wide array of scenarios, akin to the way large language models are trained and evaluated. Moreover, the unique problem setting of Embodied AI opens up the possibility for empirical evaluation to become the cornerstone of causal model assessment. Through interaction with both real and simulated environments, Embodied AI agents can generate and leverage interventional or counterfactual data, thereby permitting a more pragmatic approach to causal understanding. Thus, empirically-driven causal research under FVWMs and embodied AI is equally or more important than the theory-oriented paradigm.
@@
The principle that correlation does not imply causation is highly relevant in the field of predictive modeling. It is often misunderstood as meaning that predictive models cannot be causal. This interpretation holds some truth when models are trained solely on observational data, which are prone to confounding effects that can bias predictions about the outcomes of specific actions. However, it does not apply when a predictive model is trained with active intervention data, which allows the model to demonstrate causal properties. This aligns with the discussion in misconception I : RL models are inherently causal and can learn to take optimal actions for a variety of tasks (Kaddour et al., 2022; Richens & Everitt, 2024). Thus, causal models are inherently high-accuracy prediction models of consequences about all possible actions. Predictive models can therefore be considered causal models when trained with appropriate algorithms and datasets, even without explicit causal theory considerations. In the realm of Embodied AI, opportunities arise for gathering suitable interaction data in the physical world. This enables the integration of both causal considerations and large-scale prediction models (Ke et al., 2019; Scherrer et al., 2021; Toth et al., 2022).